OpenStudy (emilyjones284):

z1=3(cos5pi/6 +isin5pi/6) and z2=2(cos2pi/3 +isin2pi/3) Find z1z2

OpenStudy (whpalmer4):

Yikes :-)

OpenStudy (emilyjones284):

Yeah I am so confused this chapter :(

OpenStudy (whpalmer4):

\[z_1 = 3(\cos \frac{5\pi}{6} + i\sin\frac{5\pi}{6})\]\[z_2 = 2(\cos \frac{2\pi}{3} + i\sin \frac{2\pi}{3})\]Find \(z_1z_2\) Is that the correct stuff?

OpenStudy (emilyjones284):

Yeah I remember doing that in class

OpenStudy (whpalmer4):

Okay, well, how well do you know your unit circle? We'll need both sin and cos of \(\dfrac{5\pi}{6}\) and \(\dfrac{2\pi}{3}\)

OpenStudy (emilyjones284):

I know it pretty well

OpenStudy (whpalmer4):

Great. Here's your chance to prove it ;-)

OpenStudy (whpalmer4):

\[\cos \frac{5\pi}{6} =\]\[\sin \frac{5\pi}{6} =\]\[\cos \frac{2\pi}3 =\]\[\sin \frac{2\pi}3 =\]

OpenStudy (emilyjones284):

- square root 3/2 1/2 -1/2 square root of 3/2

OpenStudy (whpalmer4):

Damn, you're good :-)

OpenStudy (emilyjones284):

Haha thanks (:

OpenStudy (whpalmer4):

Okay, so now we plug those into our expressions for \(z_1, z_2\): \[z_1 = 3(\cos\frac{5\pi}{6} + i \sin\frac{5\pi}{6}) = 3( -\frac{\sqrt{3}}{2} + i\frac{1}{2})\]Can you do the same for \(z_2\), please?

OpenStudy (emilyjones284):

yeah just a second

OpenStudy (emilyjones284):

z2= 2(cos 2pi/3 + isin2pi/3) =2(-1/2 =i square root of 3/2)

OpenStudy (whpalmer4):

the = sign is really a + sign where you missed with the shift key, right? \[z_2 = 2(-\frac{1}{2}+i\frac{\sqrt{3}}{2})\]

OpenStudy (emilyjones284):

oh yeah oops haha

OpenStudy (whpalmer4):

Now we just multiply those two together, a bit messy, but hey, that's life in the algebra fast lane... \[z_1z_2 = 3( -\frac{\sqrt{3}}{2} + i\frac{1}{2})* 2(-\frac{1}{2}+i\frac{\sqrt{3}}{2})\]\[\qquad = 6( -\frac{\sqrt{3}}{2} + i\frac{1}{2})* (-\frac{1}{2}+i\frac{\sqrt{3}}{2})\] I did the easy part for you :-)

OpenStudy (emilyjones284):

Haha thanks (:

OpenStudy (whpalmer4):

Here's a hint: when I get some ugly thing like this, I look to see if there's some simple substitution I can do to make the algebra more pleasant. You've really only got two quantities here, right? \[a = \frac{\sqrt{3}}{2}\]\[b = \frac{1}{2}\]Now suddenly our problem is \[z_1z_2 = 6(-a+bi)(-b+ai)\]Multiply that out, then substitute the fractions back in and simplify...

OpenStudy (emilyjones284):

Haha alright Ill try my best

OpenStudy (emilyjones284):

So i got \[(6 \frac{ \sqrt{-3} }{ 2 } + 6 \frac{ 1 }{ 2 })\times(6\frac{ -1 }{ 2 } + i \frac{ \sqrt{3} }{ 2 })\]

OpenStudy (emilyjones284):

But I know i have to simplify it I just wanted to make sure I was right so far

OpenStudy (whpalmer4):

\[6(ab -a^2i -b^2i + abi^2)\]\[6(ab-(a^2+b^2)i -ab)\] \[6(\frac{\sqrt{3}}{4}-(\frac{3}{4}+\frac{1}{4})i - \frac{\sqrt{3}}{4}i)\]

OpenStudy (emilyjones284):

and for the last part I meant 6 square root of 3/2

OpenStudy (whpalmer4):

Oh, I wouldn't go down the road you're on, and you moved a - under the radical in the first part

OpenStudy (emilyjones284):

Jeez I knew I was doing something wrong. I currently hate math.

OpenStudy (whpalmer4):

Keep the 6 outside for as long as you can. But you might look at what I did and see if you don't like that better...

OpenStudy (whpalmer4):

Whoops, I made a typo: \[6(\frac{\sqrt{3}}{4}-(\frac{3}{4}+\frac{1}{4})i - \frac{\sqrt{3}}{4})\] Notice that the terms with the square root cancel out? \[6(-(\frac{3}{4}+\frac{1}{4})i) = -6(1i) = -6i\]

OpenStudy (emilyjones284):

Oh wow so that is the answer?

OpenStudy (anonymous):

\[z1=3e ^{\iota \frac{ 5\pi }{ 6 }},z2=2e ^{\iota \frac{ 2\pi }{ 3 }}\] \[z1z2=6e ^{\iota \left( \frac{ 5\pi }{6 }+\frac{ 2\pi }{ 3 } \right)}=6e ^{\iota \left( \frac{ 5\pi+4\pi }{6 } \right)}\] \[=6e ^{\iota \frac{ 3\pi }{2 }}=6\left( \cos \frac{ 3\pi }{2 }+\iota \sin \frac{ 3\pi }{2 } \right)=6\left( 0+\iota(-1) \right)=-6\iota \] \[\left[ \cos \frac{ 3\pi }{2 }=\cos \left( 2\pi-\frac{ \pi }{2 }\right)=\cos \frac{ \pi }{2 } =0\right]\] \[\left[ \sin \frac{ 3\pi }{2 }=\sin \left( \pi+\frac{ \pi }{2 } \right)=-\sin \frac{ \pi }{2 } =-1\right]\]

OpenStudy (whpalmer4):

Yes, \(z_1z_2 = -6i\)

OpenStudy (emilyjones284):

Well at least I understand how to do it know haha thank you (:

OpenStudy (whpalmer4):

My colleague is showing another way to do it which is possibly the one your teacher had in mind...

OpenStudy (emilyjones284):

I don't suppose you would have time to help me on about 10 other different problems would you? Haha :P

OpenStudy (whpalmer4):

Not right now, unfortunately — I've got to go run an errand for about an hour. I'll check back when I return.

OpenStudy (emilyjones284):

I actually think our teacher taught us the way that you did it

OpenStudy (emilyjones284):

Okay thanks for your help (:

OpenStudy (whpalmer4):

My way works even if the angles aren't the same, but the other way is easier, if you know it (and the angles are the same, which they usually would be).

OpenStudy (whpalmer4):

There's a cool identity that the other formula reminds me of: \[e^{\pi i} +1 = 0\]5 very special numbers, all tied together!

OpenStudy (emilyjones284):

Haha maybe I'll try to figure that way out